Bens grandfather was listening to the weather report which indicated that the maximum for the day was going to be 30 degrees celsius. He wants to know what that temperature is in farenheit using the formula celsius/ 5 = fahrenheit - 32 / 2
Help please.

Answer by rfer(16322)

F=9/5C+32
F=9/5*30+32
F=54+32
F=86